Two buses A & B leave the same bus depot , A towards north,B towards east.after four hours the distance between the two buses is 100 km .
determine the speed of the bus A if it travels at the speed of 5 km/h more than that of bus B?

Found 2 solutions by addingup, MathTherapy:
Answer by addingup(3677)

A = B + 5 per hour
A + B = 100 total (4 hours)
Per hour: 100/4 = 25
Substitute for A on a per hour basis:
B + 5 + B = 25
2B = 20
B = 10
--------------------------------------------
Check:
B: 10 x 4 = 40
A: (10 + 5) x 4 = 60
and 40 + 60 = 100 - Correct

Answer by MathTherapy(10552)


Two buses A & B leave the same bus depot , A towards north,B towards east.after four hours the distance between the two buses is 100 km .
determine the speed of the bus A if it travels at the speed of 5 km/h more than that of bus B?
Correct answer: 
This scenario involves a 3-4-5 Pythag triple, times 20.
Again, some people on here seem to have 1 aim, and that's to provide INCORRECT answers.
IGNORE those people's responses!